OpenStudy (anonymous):

the wildcats scored an average of 56 points per game during their four game season if the scores for their first three games were 48, 57 and 52what was the score of their last game?

OpenStudy (anonymous):

67

OpenStudy (anonymous):

explain please

OpenStudy (anonymous):

well, the process to finding the average, in this case 56, can be described as such: first we add up all 4 scores( we were told there were four scores) and the divide then by the number of entires, in this case 4. That means that we can reverse that process, and find out the sum of all four scores, simply by mulitplying 56 by 4 (since multiplication undoes division). So i did, 56*4 and got 224. Which means that 224 is the sum of all the scores (4 scores in total). Now, with this in mind, we are given the scores of 3 games. That is 48, 57, and 52. All we have to do now i add these three scores together,we get 157, and subtract that number from 224 (which is the sum of all 4 scores). Upon doing so, we find that the differece between 224 and 157 is 62. That means that the 4th score is 62. To check that we can simply find the average of 48,52,57 and 62.

OpenStudy (anonymous):

there are 4 games the equation of your problem is : (48 + 57 + 52 + ?)/4 = 56 (it's the formula of the average so 56* 4 - 48 - 57 - 52 = ? and so the number you want is : 67